OpenStudy (anonymous):

solving equations by graphing 3x+4y=8 x-3y=-6

OpenStudy (anonymous):

woud i add the 4y and 3y together?

OpenStudy (bahrom7893):

solve for x in the 2nd equation: x = 3y-6 Plug that into eqn above:

OpenStudy (bahrom7893):

3(3y-6)+4y = 8 9y - 18 + 4y = 8 13y = 8+18 13y = 26; y = 2

OpenStudy (bahrom7893):

y = 2 x = 3y-6 = 6-6 = 0

OpenStudy (bahrom7893):

x = 0; y = 2 Please click on become a fan if I helped! Thanks!=)

OpenStudy (bahrom7893):

sorry didnt see that it was by graphing: to solve by graphing find y in terms of x in both cases and then graph them on a calculator and ur answer will be the point where they intersect (in this case (0;2)).. I'm off to class now